# Reinforcement Learning for Monetary Policy Under Macroeconomic Uncertainty: Analyzing Tabular and Function Approximation Methods

### v1

@@ Abstract
We study how a central bank should dynamically set short-term nominal interest rates to stabilize inflation and unemployment when macroeconomic relationships are uncertain and time-varying. We model monetary policy as a sequential decision-making problem where the central bank observes macroeconomic conditions quarterly and chooses interest rate adjustments. Using publically accessible historical Federal Reserve Economic Data (FRED), we construct a linear-Gaussian transition model and implement a discrete-action Markov Decision Process with a quadratic loss reward function. We chose to compare nine different reinforcement learning style approaches against Taylor Rule and naive baselines, including tabular Q-learning variants, SARSA, Actor-Critic, Deep Q-Networks, Bayesian Q-learning with uncertainty quantification, and POMDP formulations with partial observability. Surprisingly, standard tabular Q-learning achieved the best performance (-615.13 ± 309.58 mean return), outperforming both enhanced RL methods and traditional policy rules. Our results suggest that while sophisticated RL techniques show promise for monetary policy applications, simpler approaches may be more robust in this domain, highlighting important challenges in applying modern RL to macroeconomic policy.
@@ 1 Introduction
Central banks worldwide face the same fundamental challenge of setting monetary policy under uncertainty. The traditional approach mainly relies on policy rules like the Taylor Rule taylor1993discretion, which prescribes interest rate adjustments based on inflation and output gap deviations from target. However, these rules assume known relationships between macroeconomic variables, which cannot be gaurenteed given the rapid structural changes in the global economy.
@@
This paper makes several contributions to the intersection of RL and monetary policy: (1) We implement and compare nine different RL approaches on a data-driven monetary policy environment, ranging from tabular methods to deep learning and Bayesian approaches. (2) We provide a comprehensive experimental evaluation using historical macroeconomic data from 1955-2025. (3) We demonstrate that simpler RL methods can outperform both sophisticated RL techniques and traditional policy rules in this domain. (4) We analyze the challenges and limitations of applying modern RL to modern macroeconomic policy.
@@ 2 Related Work
The application of RL to monetary policy has received growing attention in recent years. Yellen (2017) discusses the challenges facing modern central banking yellen2017goals, while Bernanke (2015) provides historical context on Federal Reserve decision-making under uncertainty bernanke2015courage .
Early work in this area focused on simple policy rules and their robustness properties taylor2016central. More recently, researchers have explored machine learning approaches to monetary policy. Nakamura and Steinsson (2018) studies the identification of monetary policy shocks and inflation targeting frameworks nakamura2018identification.
The RL literature on sequential decision-making under uncertainty is extensive sutton2018reinforcement. Tabular Q-learning remains a fundamental algorithm watkins1992q, while function approximation methods like Deep Q-Networks mnih2015human and Actor-Critic methods konda2000actor have enabled application to high-dimensional problems.
@@
Our work is different from previous studies by providing a comprehensive comparison of multiple RL approaches on a realistic monetary policy environment constructed from historical macroeconomic data. Our paper aims to deploy computer science methods, informed by established economic research, to explore how different RL models preform in monetary policy setting.
@@ 3.1 Environment and Data
We construct our state space from quarterly U.S. macroeconomic data obtained from the Federal Reserve Economic Data (FRED) API, spanning 1955-2025. The state vector comprises four key variables: inflation rate $\pi_{t}$ (year-over-year CPI), unemployment rate $u_{t}$, output gap $y_{t}=\frac{\text{GDP}_{t}-\text{GDP}^{\text{pot}}_{t}}{\text{GDP}^{\text{pot}}_{t}}\times 100$, and policy rate $i_{t}$ (federal funds rate). These variables help capture the state of the economy.
@@ 3.3 Baselines and Evaluation
We compare against standard monetary policy benchmarks: (1) Taylor Rule: $i^{*}_{t}=r^{*}+\pi_{t}+\phi_{\pi}(\pi_{t}-\pi^{*})+\phi_{y}y_{t}$ with $\phi_{\pi}=0.5,\phi_{y}=0.5$. (2) Taylor Hyperparameter Tuned: Literature-calibrated version with $\phi_{\pi}=1.5,\phi_{y}=0.5$ for stronger inflation response that match aforementioned contemporary research. (3) Naive Hold: Maintains current interest rate (no change, chooses hold every time).
@@ 5.1 Limitations
The reward function, while based on established dual mandate principles, involves somewhat arbitrary weight choices. Alternative situational specifications might favor different algorithmic approaches. Additionally, our historical data period includes various economic regimes (Great Moderation, Financial Crisis, etc.) that may not generalize to future economic conditions. Constructing the different models for different types of economic environment using more granular data could enhance future findings.
Lastly, adding more complexity to our problem statement could demonstrate an interesting trade-off between computational efficiency and accuracy of results. For instance, future studies could incorporate more economic indicators such as stock market indexes and mortgage rates.
@@ 6 Conclusion
We conducted a comprehensive comparative evaluation of various reinforcement learning approaches for monetary policy under macroeconomic uncertainty. Using historical Federal Reserve data from FRED and a realistic simulation environment, we compared nine RL methods against more traditional macroeconomic policy baselines like Taylor’s and Philip’s laws.
@@
Author Contributions
We all contributed equally to this project. We worked collaboratively on the design of the research framework, development of the simulation environment, implementation of the reinforcement learning methods, execution of experiments, analysis of results, and writing of the paper.
@@ Acknowledgments
We thank the CS238 staff for guidance throughout this project. We acknowledge the Federal Reserve Economic Data (FRED) system for providing the data used in our experiments.

### v2

@@ Abstract
We study how a central bank should dynamically set short-term nominal interest rates to stabilize inflation and unemployment when macroeconomic relationships are uncertain and time-varying. We model monetary policy as a sequential decision-making problem where the central bank observes macroeconomic conditions quarterly and chooses interest rate adjustments. Using publicly accessible historical Federal Reserve Economic Data (FRED), we construct a linear-Gaussian transition model and implement a discrete-action Markov Decision Process with a quadratic loss reward function. We chose to compare nine different reinforcement learning style approaches against Taylor Rule and naive baselines, including tabular Q-learning variants, SARSA, Actor-Critic, Deep Q-Networks, Bayesian Q-learning with uncertainty quantification, and POMDP formulations with partial observability. Notably, despite its simplicity, standard tabular Q-learning achieved the best performance (-615.13 ± 309.58 mean return), outperforming both enhanced RL methods and traditional policy rules. Our results suggest that while sophisticated RL techniques show promise for monetary policy applications, simpler approaches may be more robust in this domain, highlighting important challenges in applying modern RL to macroeconomic policy.
@@ 1 Introduction
Central banks worldwide face the same fundamental challenge of setting monetary policy under uncertainty. The traditional approach mainly relies on policy rules like the Taylor Rule taylor1993discretion, which prescribes interest rate adjustments based on inflation and output gap deviations from target. However, these rules assume known relationships between macroeconomic variables, which cannot be guaranteed given the rapid structural changes in the global economy.
@@
This paper makes several contributions to the intersection of RL and monetary policy: (1) We implement and compare nine different RL approaches on a data-driven monetary policy environment, ranging from tabular methods to deep learning and Bayesian approaches. (2) We provide a comprehensive experimental evaluation using historical macroeconomic data from 1955-2025. (3) We demonstrate that simpler RL methods can outperform both sophisticated RL techniques and traditional policy rules in this domain. (4) We analyze the challenges and limitations of applying RL to macroeconomic policy.
@@ 2 Related Work
The application of RL to monetary policy has received growing attention in recent years. Yellen (2017) discusses the challenges facing modern central banking, especially regarding uncertainty about how evolving economic and productivity trends complicate policy, motivating frameworks that can systematically adapt to shifting macroeconomic relationships yellen2017goals. Bernanke (2015) provides historical context on Federal Reserve decision-making under uncertainty, demonstrating the limits of conventional tools like interest-rate cuts during the 2007–2008 financial crisis, reinforcing the notion that effective policy must contend with unforeseeable shocks and feedback effects, features that machine learning and reinforcement learning frameworks explicitly seek to model and optimize under uncertainty bernanke2015courage .
Early work in this area focused on simple policy rules and their robustness properties taylor2016central. More recently, researchers have explored machine learning approaches to monetary policy. Nakamura and Steinsson (2018) studied the identification of monetary policy shocks and inflation targeting frameworks nakamura2018identification.
The RL literature on sequential decision-making under uncertainty is extensive sutton2018reinforcement. Tabular Q-learning remains a fundamental algorithm for low-dimensional, discretized environments watkins1992q, while function approximation methods like Deep Q-Networks mnih2015human and Actor-Critic methods konda2000actor have enabled application to high-dimensional problems, albeit with increased instability and sensitivity to hyperparameters.
@@
Our work is different from previous studies by providing a comprehensive comparison of multiple RL approaches on a realistic monetary policy environment constructed from historical macroeconomic data. Our paper aims to deploy computer science methods, informed by established economic research, to explore how different RL models perform in monetary policy setting.
@@ 3.1 Environment and Data
We construct our state space from quarterly U.S. macroeconomic data obtained from the Federal Reserve Economic Data (FRED) API fred_api_docs, spanning 1955-2025. The state vector comprises four key variables: inflation rate $\pi_{t}$ (year-over-year CPI), unemployment rate $u_{t}$, output gap $y_{t}=\frac{\text{GDP}_{t}-\text{GDP}^{\text{pot}}_{t}}{\text{GDP}^{\text{pot}}_{t}}\times 100$, and policy rate $i_{t}$ (federal funds rate). These variables help capture the state of the economy.
@@ 3.3 Baselines and Evaluation
We compare against standard monetary policy benchmarks: (1) Taylor Rule: $i^{*}_{t}=r^{*}+\pi_{t}+\phi_{\pi}(\pi_{t}-\pi^{*})+\phi_{y}y_{t}$ with $\phi_{\pi}=0.5,\phi_{y}=0.5$. (2) Taylor Hyperparameter Tuned: Literature-calibrated version with $\phi_{\pi}=1.5,\phi_{y}=0.5$ for stronger inflation response that matches aforementioned contemporary research. (3) Naive Hold: Maintains current interest rate (no change, chooses hold every time).
@@ 5.1 Limitations
The reward function, while based on established dual mandate principles, involves somewhat arbitrary weight choices. Alternative situational specifications might favor different algorithmic approaches. Additionally, our historical data period includes various economic regimes (Great Moderation, Financial Crisis, etc.) that may not generalize to future economic conditions. Constructing different models for different types of economic environment using more granular data could enhance future findings.
One central limitation is that agents interact with a fixed transition model estimated from historical data, not the real macroeconomic system. While enabling controlled experimentation, this means our models optimize within a single consistent environment rather than adapting to genuinely changing dynamics. The observed performance thus reflects suitability for simplified, stationary control problems, not robustness under true macroeconomic nonstationarity.
Relatedly, tabular Q-learning’s success may stem from strong inductive biases: manual discretization and low-dimensional representations impose structured abstraction aligned with linear dynamics, reducing variance and stabilizing learning. Function approximation methods may be disadvantaged by unnecessary flexibility. Our findings suggest careful state abstraction can dominate algorithmic sophistication in stylized macroeconomic settings, rather than endorsing simpler algorithms generally.
Finally, adding more complexity to our problem statement could demonstrate an interesting trade-off between computational efficiency and accuracy of results. For instance, future studies could incorporate more economic indicators such as stock market indexes and mortgage rates.
@@ 6 Conclusion
We conducted a comprehensive comparative evaluation of various reinforcement learning approaches for monetary policy under macroeconomic uncertainty. Using historical Federal Reserve data from FRED and a realistic simulation environment, we compared nine RL methods against more traditional macroeconomic policy baselines like Taylor’s and the Phillips curve.
@@
Simultaneously, our results highlight an important methodological caveat for applying reinforcement learning to economic policy analysis: strong empirical performance in a simulated, data-fitted environment does not necessarily translate to robustness in real-world policymaking contexts. Because both training and evaluation occur within the same estimated transition model, differences across algorithms primarily capture optimization efficiency under known dynamics, rather than resilience to structural breaks, policy-induced feedback loops, or rare macroeconomic events.
@@ Acknowledgments
We thank Professor Mykel J. Kochenderfer and the Stanford AA228/CS238 staff for guidance throughout this research. We thank the Federal Reserve Economic Data (FRED) system for providing the data used in our experiments.